Options:
A.
B.
C.
D.
Sign in to view the correct answer and detailed explanation.
Ulearngo is an independent study app. It is not affiliated with, endorsed by, or representing West African Examinations Council (WAEC). Always confirm official exam, admission, registration, syllabus, timetable, and result information from WAEC official website, WAEC Nigeria.
Options:
1 older comment kept for context.
Victor Titus
P(at least one head) = 1 - P(4 tails)
Let \(p = \frac{1}{2}\) = probability of head and \(q = \frac{1}{2}\) = probability of tail.
\((p + q)^{4} = p^{4} + 4p^{3}q + 6p^{2}q^{2} + 4pq^{3} + q^{4}\)
P(4 tails) = \(q^{4} = (\frac{1}{2})^{4} = \frac{1}{16}\)
P(at least one head) = \(1 - \frac{1}{16} = \frac{15}{16}\)